OpenStudy (anonymous):

use lim = f(h+a)-f(a)/h to find the derivative h->0 f(x)=2x^3 a=10

OpenStudy (anonymous):

I got 6000 but that doesn't seem right

OpenStudy (anonymous):

can you post what you did ?

OpenStudy (anonymous):

2(10+h)^3-8000/h= 8000+8h^3-2000/h

OpenStudy (anonymous):

the h cancels the cube and makes it a square 8000+8h^2-2000 and since h is going to 0 than 8h^2 is 0 and when you combine 8000-2000 you get 6000

OpenStudy (anonymous):

I got 600

OpenStudy (anonymous):

isn't 10*10*10= 1000

OpenStudy (anonymous):

I'll post it - I might have a mistake.

OpenStudy (anonymous):

\[\frac{1}{h}*(2(h+a)^3-2a^3)\] \[=\frac{1}{h}*(\cancel{2a^3}+6a^2h+6ah^2+h^3-\cancel{2a^3})\] \[=6a^2+6ah+h^2\] \[\lim_{x \rightarrow 0} 6a^2+6ah+h^2=6a^2=6*10^2=600\]

OpenStudy (anonymous):

in the lim above it should be h->0 not x->0 obviously :)

OpenStudy (anonymous):

alright thanks